OpenStudy (anonymous):

√6*√14= √6=2.449 √14=3.741 9.161709 Is this correct?

OpenStudy (anonymous):

Yes but I'm not sure that's what the question wants. Probably wants to see you manipulate the radicals the simplify

OpenStudy (anonymous):

Manipulate the radicals how?

OpenStudy (anonymous):

3√2*7√2??

OpenStudy (anonymous):

Like sqrt(6) * sqrt(14) = sqrt(6 * 14) = sqrt(84) = sqrt(4 * 21) = 2sqrt(21) Careful, when you take something out of a radical. See how the 4 comes out as a 2 Thats because sqrt(4 * 21) = sqrt(4) * sqrt(21) = 2sqrt(21)

OpenStudy (anonymous):

ahhhhhhh I see said the blind man.

OpenStudy (anonymous):

@Xavier Thanks